I am supposed to solve the equation 3g^{2}-24g+27=0 by completing the square and round your solutions to the nearest hundredth if necessary.

I came up with the answers approximately 23.62 and approximately 0.38  

However,, the textbook answers are 1.35 and 6.65

How did they get these answers?

Please show complete worked-out solutions.

A sincere thank you!
Answer #1

Ans we have given = 3 3g²- 249 +27=0 3 [92-89 +3=0 3[02 x29 +(3? - (+)= 0 3C9289 +16-76 +95 = 0 3 [92-39+16 - 7) = 0 3 [82-89